Find the lenghth of a rectangle when the width is 2 inches shorter and the perimeter is 80. Set up the equation for the perimeter involving only (L)
and then sove the equation for the length and the width.

Answer by checkley77(12844)

W=L-2
2L+2W=80
2L+2(L-2)=80
2L+2L-4=80
4L=80+4
4L=84
L=84/4
L=21 ANSWER FOR THE LENGTH.
W=21-2
W=19 ANSWER FOR THE WIDTH.
PROOF:
2*21+2*19=80
42+38=80
80=80